OpenStudy (cathyangs):

If a is not equal to b, and a^3-b^3=19x^3, and a-b=x, which of the following conclusions is correct? a=3x a=3x or a=-2x a=-3x or a=2x a=3x or a=2x a=2x

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2+ab+b^2)=19x^3\] \[a-b=x\] \[a^2+ab+b^2=19x^2\]

OpenStudy (anonymous):

i guess that does not solve the problem does it?

OpenStudy (cathyangs):

I got that far...but I don't know how to go from there.

OpenStudy (anonymous):

i don't know lets see what we can try \[a-b=x\] \[(a-b)^2=x^2=a^2-ab+b^2\] so we have \[x^2=a^2-ab+b^2\] and \[19x^2=a^2+ab+b^2\]

OpenStudy (anonymous):

sorry it should be \[a^2-2ab+b^2=x^2\]

OpenStudy (anonymous):

19x^2-x^2=a^2+ab+b^2-(a^2-2ab+b^2)\] \[18x^2=-3ab\] well that still doesn't do it does it?

OpenStudy (anonymous):

i can't seem to get rid of the "b"

OpenStudy (zarkon):

try some of the a's given...a=3x or a=-2x work

OpenStudy (precal):

if a=3x then b=2x Is there a typo in the solutions given?

OpenStudy (mertsj):

\[a^+ab+b^2=19x^2\] \[b=a-x\] \[a^2+a(a-x)+a^2-2ax+x^2=19x^2\] \[a^2+a^2-ax+a^2-2ax+x^2=19x^2\] 3a^2-3ax=18x^2

OpenStudy (mertsj):

\[3a^2-3ax-18x^2=0\]

OpenStudy (mertsj):

\[a^2-ax-6x^2=0\]

OpenStudy (precal):

@cathyangs Is there a typo in the solutions given?

OpenStudy (mertsj):

\[(a-3x)(a+2x)=0\]

OpenStudy (anonymous):

very nice

OpenStudy (mertsj):

\[a=3x or -2x\]

OpenStudy (mertsj):

Thank you. That really means something coming from The Great One as it were.

OpenStudy (anonymous):

jackie gleason was the great one!

OpenStudy (mertsj):

So you're in excellent company.